Find the total capacitance for three capacitors connected in series, given their 

individual capacitances are 1.000, 5.000, and 8.000 µF. (0.755µF)



1
Expert's answer
2022-03-15T10:33:53-0400
"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C}=\frac{1}{1.000\:\rm \mu F}+\frac{1}{5.000\:\rm \mu F}+\frac{1}{8.000\:\rm \mu F}=1.325"

"C=0.755\:\rm \mu F"